OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (anonymous):

Since it is perpendicular to line y=-x-3, its slope must be 1. (ab=-1 for two lines with slope a and b to be perpendicular) So, the equation would be y=x+a (a is a constant) Now plug in (2,3) 3=2+a a=1 y=x+1

OpenStudy (amorfide):

perpendicular to the line y=-x-3 means the gradient of the perpendicular line is the negative reciprocal of y=-x-3 so the gradient = -1/-1=1